HETAV PATEL

Grade 11,

The electron in ahydrogen atom makes transition fromM toL shell. The ratio of magnitude of centripetal acceleration of the electron

The electron in ahydrogen atom makes transition fromM toL shell. The ratio of magnitude of centripetal acceleration of the electron
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
As we know
acceleration is proportional to v^2/r
 
 
a = v^2 /r = 1/ n^4
 
hence a1 / a2 = 2^4 / 3^4 = 16 / 81